+ IA()
media%2F175%2F17534a05-193b-4222-9755-3f
media%2F085%2F085f4c14-0862-41c3-bb00-74
Answer #1

1. 2 Al + 6 HNO3 === 2 Al(NO3)3 + 3 H2 , this is a single displacement reaction the Aluminum displaces the hydrogen.

2. Oxalic acid 3 M, you used 11.6 ml or 0.0116 L

moles of oxalic acid = Molarity * volume = 3 * 0.0116 = 0.0348 moles

you need 2 moles of KOH for every mole of oxalic acid so if you use 0.0348 moles of oxalic acid , then you had

2 * 0.0348 = 0.0696 moles of KOH

Molarity = moles / volume = 0.0696 / 0.025 = 2.784 M ; the 0.025 comes from the 25 ml mentioned in the statement.